Which of the options in this scenario is INCORRECT?

A pet store has 8 cats, 12 dogs and 4 rabbits. The ratio 1:3 compares:


1
Expert's answer
2021-04-27T08:56:30-0400

Given, A pet store has 8 cats, 12 dogs and 4 rabbits.

Ratio of cats and dogs,

"\frac{8}{12}=\frac{2}{3}"

Ratio of rabbits and cats,

"\frac{4}{8}=\frac{1}{2}"

Ratio of rabbits and dogs,

"\frac{4}{12}=\frac{1}{3}"

Since, given ratio is 1:3.

Therefore, ratio of cats and rabbits, ratio of rabbits and dogs are incorrect.